how to prepare 50ml solution of 5%nacl
1
Expert's answer
2019-05-01T04:28:43-0400

%=m(solute)/(m(solute)+m(solvent) x 100

Lets assume waster as a solvent. Its density is 1 g/ml. Therefore m=V (1g=1 ml).

Density of 5% NaCl solution is 1.07 g/ml.

m(NaCl5%)=50x1.07=53.5 g

m(NaCl)=0.05x53.5=2.7 g

m(H2O)=53.5-2.7=50.8 g

To prepare 50 ml solution of 5% NaCl take 50.8 g of H2O and add 2.7 g of NaCl (100%).